OpenStudy (anonymous):

Carla and Benicio work in a men's clothing store. They earn commission from each suit and each pair of shoes they sell. For selling 3 suits and one pair of shoes, Carla has earned $47 in commission. For selling 7 suits and 2 pairs of shoes, Benicio has earned $107 in commission. How much do the salespeople earn for the sale of a suit?...for the sale of a pair of shoes?

OpenStudy (anonymous):

3x+y=47 7x+2y=107 y=47-3x 7x+94-6x=107 x=13 y=8

OpenStudy (anonymous):

confused? x is suits, y is shoes